OpenStudy (anonymous):

Equation of circle in standard form with the following center and radius: Center (-8, -3), radius = 2

OpenStudy (anonymous):

(x+8)^2+(y+3)^2=2^2

OpenStudy (anonymous):

Ooops I mean radius 2 (3 squared)

OpenStudy (anonymous):

Equation of circle in standard form is (x-h)^2 + (y-k)^2 =r^2 now your center is in point (h,k)=(-8,-3),,just plug in your given point to the formula (x-h)^2 + (y-k)^2 =r^2 (x-(-8))^2 + (y-(-3))^2 =2^2 (x+8)^2+(y+3)^2=2^2 or (x+8)^2+(y+3)^2=4 ans...

OpenStudy (anonymous):

did you get it wow? or is it clear now?..lol